OpenStudy (anonymous):

RSV has coordinates R(2, 1), S(3, 2), and V(2, 6). A translation maps points R to R' at (–4, 8). What are the coordinates for S' for this translation?

OpenStudy (accessdenied):

What kind of transformation is a translation? When we take the figure, what do we 'do' with it to make a translation?

OpenStudy (accessdenied):

hmm, cant tell if you're not here, so ill just continue. A translation moves the figure without resizing it or rotating it. So, when we translate it, we are moving the x-values by a constant and the y-values by a constant. From R to R', R(2,1) -> R'(-4,8). If we wanted to define the shift, we'd use something like (x + h, y + k). (x,y) is R, (x+h,y+k) is R'. 2 + h = -4, 1 + k = 8 h = -6, k = 7 The shift is defined for (x,y); (x-6, y+7). We can use (x,y) = S(3,2) to find S'.

OpenStudy (accessdenied):

here's a visual of it